# BoundMatch: Boundary detection applied to semi-supervised segmentation for urban-driving scenes

### v1

@@ 4.3 Ablation Studies and Insights
In this section, we conduct ablation studies to analyze the impact of each proposed component and design choice. We report results on the Cityscapes dataset under the $\nicefrac{{1}}{{16}}$ labeled data split (unless explicitly stated), and our evaluation metrics include the mean Intersection-over-Union (IoU), boundary IoU (BIoU), and boundary F1 score (BF1).
@@ 4.3.1 Individual efficacy of the proposed components
Table 5: Component Analysis.
@@
75.06 54.93 59.50
75.91 55.25 60.22
76.21 55.95 61.97
76.10 55.70 61.64
76.38 55.97 62.46
Tab. 5 reports the quantitative results for various combinations of our proposed components, starting from the SAMTH baseline which lacks explicit boundary supervision. The addition of our core BCRM framework alone results in notable gains in both segmentation and boundary quality metrics. Further incorporating the BSF module builds upon this, enhancing performance by effectively integrating learned boundary cues into the segmentation head. Finally, adding the SGF module provides complementary benefits; its strategy of refining boundary predictions using spatial gradients derived from the segmentation mask leads to the highest overall performance observed in our ablation.
@@ 4.3.2 Binary vs Semantic Boundaries
+Multi-Label 76.38 55.97 62.46
An important design decision in our method is whether to model boundaries as binary or multi-label (semantic). As shown in Tab. 6, while both formulations improve upon the SAMTH baseline, the semantic boundary formulation achieves the highest improvement in segmentation performance (with IoU increasing from 75.06% to 76.38%). Performance gains from semantic boundaries are supported by experiments in the fully supervised setting for urban scene segmentation [11]. We believe this forces the model to learn not just geometric structures but also semantic context of the boundaries. This suggests that incorporating semantic context into the boundary predictions allows the network to better differentiate between object classes at the boundaries.
@@ 4.3.3 Boundary Evaluation
Qualitative evaluations of the boundary predictions are presented in Fig. 6. We observe that the proposed multi-label boundary formulation yields sharper and more precise object contours compared to the baseline and binary boundary formulations. These qualitative results corroborate the quantitative improvements reported in the previous subsections.
While mIoU remains the de facto standard for assessing overall segmentation quality, it is well known to be insensitive to errors occurring specifically at object contours [10, 59]. To explicitly quantify boundary fidelity, we adopt two complementary metrics: Boundary IoU (BIoU) and Boundary F1 Score (BF1).
Boundary IoU (BIoU):
$\mathrm{BIoU}\;=\;\frac{1}{C}\sum_{c=1}^{C}\frac{\sum_{n}\sum_{(i,j)\in BD(y_{n})}\bigl{[}\hat{y}_{n,c,i,j}\,\wedge\,y_{n,c,i,j}\bigr{]}}{\sum_{n}\sum_{(i,j)\in BD(y_{n})}\bigl{[}\hat{y}_{n,c,i,j}\,\vee\,y_{n,c,i,j}\bigr{]}},$ (14)
where $BD(y)=y\oplus\mathrm{MinPool}(y)$ denotes the binary boundary mask obtained by applying a 3×3 min-pooling (stride 1) to the ground-truth segmentation map $y$ and taking the pixel-wise XOR ($\oplus$) with the original mask, and $\wedge$ and $\vee$ denote pixel-wise logical AND and OR, respectively [60]. Because it ignores interior pixels, BIoU directly measures contour alignment and penalizes both over- and under-segmentation along object edges. It is therefore particularly sensitive to boundary precision but agnostic to interior region accuracy. We use a 5 pixel radius for the boundary mask in our experiments.
Boundary F1 Score (BF1):
[…]
$\mathrm{Prec}_{c}\;=\;\frac{\sum_{n,h,w}\bigl{[}\mathrm{BD}(\hat{y}_{n})_{c,h,w}\,\wedge\,\mathrm{MaxPool}\bigl{(}\mathrm{BD}(y_{n})\bigr{)}_{c,h,w}\bigr{]}}{\sum_{n,h,w}\mathrm{BD}(\hat{y}_{n})_{c,h,w}},$ (16)
and
$\mathrm{Rec}_{c}\;=\;\frac{\sum_{n,h,w}\bigl{[}\mathrm{MaxPool}\bigl{(}\mathrm{BD}(\hat{y}_{n})\bigr{)}_{c,h,w}\,\wedge\,\mathrm{BD}(y_{n})_{c,h,w}\bigr{]}}{\sum_{n,h,w}\mathrm{BD}(y_{n})_{c,h,w}}.$ (17)
BF1 thus provides insight into the trade-off between missing fine boundary details (low recall) and producing spurious edges (low precision), while allowing for minor spatial misalignments. We use a 5 pixel tolerance radius for our experiments.
Complementary Insights:
BIoU emphasizes exact contour overlap, making it well suited for tasks requiring crisp, well-aligned boundaries. BF1, by explicitly balancing precision and recall under a spatial tolerance, is more robust to slight shifts but sensitive to fragmented or noisy boundary predictions. Reporting both metrics alongside mIoU therefore provides a holistic view of a model’s segmentation performance, particularly its ability to delineate object contours accurately.
Analysis of BIoU and BF1 Scores:
Tab. 5 and Tab. 6 reports the BIoU and BF1 scores for various configurations of our proposed components and boundary types. We observe that the addition of boundary-aware components (BCRM, BSF, and SGF) consistently improves both BIoU and BF1 scores, with BSF and SGF having similar gains. Combining the two components results in a significant boost compared to BCRM, with the highest boundary metrics. We also found that Binary boundaries are less effective than Multi-Label boundaries, which is consistent with the segmentation results. These improvements in BIoU and BF1 underscore the effectiveness of our approach in capturing fine-grained boundary details. The refined boundaries contribute not only to visually sharper segmentation maps but also to more accurate delineation of object contours, which is critical for applications requiring precise localization.
@@ 4.3.4 Comparisons with BoundaryMatch
Table 7: Comparison between our BoundMatch method and BoundaryMatch. Performance values for UniMatch [9] and BoundaryMatch [15] are included for context. For BoundaryMatch, we also reimplemented their method and trained the method with the same setting for fairness. All methods use ResNet-101 as the backbone.
Methods VOC (92) City. ($\nicefrac{{1}}{{16}}$)
UniMatch 75.2 76.6
@@
BoundaryMatch (reproduced) 75.5 76.8
SAMTH+BoundMatch 76.6 77.9
UniMatch+BoundMatch 76.0 77.4
We compared our BoundMatch approach against BoundaryMatch [15], a recent method that also utilizes boundary CR for SS-SS. For a fair comparison based on our experimental setup, we reimplemented BoundaryMatch and trained it using identical settings. Quantitative results are presented in Tab. 7.
While BoundaryMatch shows minor gains over its UniMatch baseline, our SAMTH+BoundMatch method significantly outperforms the reproduced BoundaryMatch implementation by 1.1% mIoU on both Pascal VOC (92 split) and Cityscapes ($\nicefrac{{1}}{{16}}$ split). Furthermore, integrating BoundMatch’s principles with UniMatch (UniMatch+BoundMatch) also surpasses BoundaryMatch’s performance on both datasets, although the improvement margin is larger when using our SAMTH baseline, suggesting potential synergy effects.
@@ 4.3.5 Effect of Harmonious BN update strategy
Table 8: Effect of Harmonious BN Update.
Methods VOC (92) City. ($\nicefrac{{1}}{{16}}$)
@@
SAMTH 70.7 +7.9 75.1 +5.7
Figure 7: Training curves of AugSeg, CutMix-MT, and SAMTH. The bold lines are method trained with HBN update strategy, while the dotted lines are ones without.
To further examine the effect of our Harmonious Batch Normalization (HBN) update strategy, we compare different BN update approaches in Tab. 8. Methods like CutMix-MT, ReCo, and U2PL also update the teacher’s BN statistics via forward passes. However, unlike HBN (which processes the complete batch through the teacher specifically for BN updates, as described in Sec. 3.3), these methods often introduce discrepancies between the inputs used for the teacher’s BN statistics and the student’s training inputs. Other approaches, such as iMAS and AugSeg, update the teacher’s BN buffers via EMA based on the student’s statistics. As demonstrated in the table, incorporating our HBN strategy consistently yields significant improvements in segmentation performance across various base methods. For example, applying HBN to SAMTH results in a 5.7% absolute improvement in mean IoU on Cityscapes (from 69.4% to 75.1%). The training curves in Fig. 7 provide further evidence, indicating that HBN stabilizes the training process and accelerates convergence, thus highlighting its practical benefits.
@@ 4.3.6 Computational Cost
Table 9: Computation Cost of methods. We report the average iteration time in seconds and the GPU memory requirements of a single GPU during training. We also report FLOPs, number of parameters, and FPS during inference.
Methods IoU Time Memory FLOPs Params FPS
UniMatch 75.3 1.21s 20.6G 191G 40.5M 23.4
[…]
+BoundMatch 76.4 1.59s 16.5G 192G 40.6M 19.0
Tab. 9 summarizes the computational cost of our approach in terms of inference time and memory consumption. Although SAMTH+BoundMatch incurs a higher computational overhead compared to the baseline UniMatch and SAMTH (with inference time increasing from 0.92 s to 1.59 s and memory usage from 9.0 G to 16.5 G), the performance gains in segmentation accuracy (with IoU increasing to 76.38%) justify the additional cost. The memory usage increases in the boundary detection head due to convolution of spatially large feature maps from the early stages of the backbone. This trade-off is especially acceptable in scenarios where segmentation precision is of paramount importance. Furthermore, the computational cost of SAMTH+BoundMatch is still comparable to that of UniMatch, while achieving better performance.
@@ 4.4 More Real-World SS-SS Setting
Inspired by the evaluation protocol in [47], we assess SS-SS in a realistic scenario: using all 2,975 Cityscapes training images as labeled data and the additional 19,997 “extra” images as truly unlabeled data. In this setting, the fully supervised baseline is already strong, making further gains challenging yet highly relevant for practical deployment. As shown in Tab. 10, SAMTH+BoundMatch achieves consistent improvements over both SAMTH and UniMatch, particularly in boundary-focused metrics (BIoU and BF1), demonstrating its effectiveness at leveraging large unlabeled pools in real-world applications.
@@ 4.5 Application to lightweight models
Method IoU BIoU BF1 FLOPs Params FPS
DLV3+ MV2 73.4 53.9 56.5 89.7G 5.8M 36.0
@@
We further evaluate BoundMatch on lightweight segmentation architectures suitable for resource-constrained devices. Specifically, we apply our method to a MobileNet-V2 backbone [62] with DeepLabV3+ and to AFFormer-Tiny, a compact Transformer model [22]. For the latter, we compare against Mobile-Seed [12], a strong baseline for semi-supervised segmentation on AFFormer-Tiny.
@@
Notably, when applied to AFFormer-Tiny, BoundMatch not only surpasses the segmentation quality of the specialized Mobile-Seed method but also achieves higher throughput (FPS) with slightly lower FLOPs. This highlights the efficiency benefits of our MTL architecture combined with SS-SS training on additional unlabeled data, even compared to methods specifically tailored for lightweight semi-supervised learning.

### v2

@@ 4.3 Ablation Studies and Insights
In this section, we conduct ablation studies to analyze the impact of each proposed component and design choice. We report results on the Cityscapes dataset under the $\nicefrac{{1}}{{16}}$ labeled data split (unless explicitly stated), and our evaluation metrics include the mean Intersection-over-Union (IoU), boundary IoU (BIoU), and boundary F1 score (BF1). We also report comparisons with BoundaryMatch, computational costs, and BoundMatch’s effectiveness in real-world semi-supervised learning settings.
@@ 4.3.1 Individual efficacy of the proposed components
Table 6: Component analysis of our framework.
Modules enabled Metrics (%)
@@
$75.05$ $54.75$ $59.28$
$75.91$ $55.25$ $60.22$
$76.21$ $55.95$ $61.97$
$76.10$ $55.70$ $61.64$
76.49 56.09 62.72
Tab. 6 reports the quantitative results for various combinations of our proposed components, starting from the SAMTH baseline which lacks explicit boundary supervision. The addition of our core BCRM framework alone results in notable gains in both segmentation and boundary quality metrics. Further incorporating the BSF module builds upon this, enhancing performance by effectively integrating learned boundary cues into the segmentation head. Finally, adding the SGF module provides complementary benefits; its strategy of refining boundary predictions using spatial gradients derived from the segmentation mask leads to the highest overall performance observed in our ablation. Per-class performance between SAMTH and SAMTH+BoundMatch is provided in Appendix D.
4.3.2 Analysis of Hyperparameters
(a) $\tau_{bdry}$
(b) $\lambda_{bdry}$
Figure 6: Analysis of the hyperparameters introduced by BoundMatch. (a) Varying the boundary threshold $\tau_{bdry}$; (b) Varying the boundary loss weight $\lambda_{bdry}$. For (b), we set $\tau_{bdry}=0.5$. The results are reported on the Cityscapes dataset with $\nicefrac{{1}}{{16}}$ and $\nicefrac{{1}}{{8}}$ split.
Fig. 6 presents the impact of two key hyperparameters introduced by BoundMatch: the boundary threshold $\tau_{bdry}$ and boundary loss weight $\lambda_{bdry}$. For the boundary threshold, we find that $\tau_{bdry}=0.5$ yields optimal performance across both data splits. This moderate threshold balances two factors: boundary predictions typically have low probability values due to the sparsity of boundary pixels, yet setting the threshold too high would exclude valuable boundary information from the consistency regularization process.
Regarding the boundary loss weight $\lambda_{bdry}$, optimal performance is achieved with values around 1.0 for both splits. Interestingly, the $\nicefrac{{1}}{{8}}$ split shows greater robustness to variations in $\lambda_{bdry}$ compared to the $\nicefrac{{1}}{{16}}$ split, suggesting that increased labeled data provides more stable training dynamics that are less sensitive to the precise weighting of the boundary loss.
@@ 4.3.3 Binary vs Semantic Boundaries
+ Multi-Label 76.49 56.09 62.72
An important design decision in our method is whether to model boundaries as binary or multi-label (semantic). As shown in Tab. 7, while both formulations improve upon the SAMTH baseline, the semantic boundary formulation achieves the highest improvement in segmentation performance (with IoU increasing from 75.05% to 76.49%). Performance gains from semantic boundaries are supported by experiments in the fully supervised setting for urban scene segmentation [11]. We believe this forces the model to learn not just geometric structures but also semantic context of the boundaries. This suggests that incorporating semantic context into the boundary predictions allows the network to better differentiate between object classes at the boundaries.
@@ 4.3.4 Boundary Evaluation
Fig. 7 presents qualitative boundary predictions at different stages of our framework. The BCRM component alone produces thicker boundaries with more artifacts, while the full BoundMatch framework generates significantly sharper and cleaner boundaries. This visual improvement demonstrates the crucial role of the BSF and SGF modules in refining boundary predictions through their fusion mechanisms.
To quantitatively assess these boundary improvements, we employ two specialized metrics beyond standard mIoU, which is known to be insensitive to boundary-specific errors [10, 63]. We adopt Boundary IoU (BIoU) and Boundary F1 Score (BF1), which explicitly measure boundary alignment and detection quality, respectively (see Appendix C for detailed definitions).
The quantitative results in Tab. 6 and Tab. 7 confirm our qualitative observations. Each boundary-aware component contributes to improved boundary metrics: BCRM alone increases BF1 from 59.28% to 60.22%, while BSF and SGF provide additional gains of approximately 1.7% and 1.4% respectively. Crucially, combining all components yields the highest performance (62.72% BF1), demonstrating their complementary nature. Furthermore, semantic boundaries consistently outperform binary boundaries across all metrics, validating our design choice to use class-specific boundary supervision. These boundary quality improvements directly translate to better segmentation accuracy, particularly for safety-critical applications requiring precise object delineation. Per-class analysis is provided in Appendix D.
@@ 4.3.5 Comparisons with BoundaryMatch
Table 8: Comparison between our BoundMatch and BoundaryMatch [15]. BoundaryMatch is our reproduced results. All models use DeepLabV3$+$ with ResNet-101.
Methods VOC (92) Cityscapes ($\nicefrac{{1}}{{16}}$)
IoU IoU BIoU BF1
@@
UniMatch $75.2$ $76.6$ $57.0$ $63.4$
BoundaryMatch $75.5$ $76.5$ $56.7$ $63.5$
SAMTH + BoundMatch (ours) 76.6 77.9 58.5 65.3
UniMatch + BoundMatch (ours) $76.0$ $77.4$ $58.0$ $64.5$
We compared BoundMatch against BoundaryMatch [15], a recent method that also utilizes boundary CR for SS-SS. For a fair comparison based on our experimental setup, we reimplemented BoundaryMatch and trained it using identical settings. Quantitative results are presented in Tab. 8 which also includes the reported mIoU for BoundaryMatch.
BoundaryMatch achieves higher mIoU compared to UniMatch baseline in Pascal VOC, but does not seem to improve IoU on Cityscapes. UniMatch+BoundMatch achieves +0.9% mIoU and SAMTH+BoundMatch achieves an even higher +1.4% compared to BoundaryMatch on Ciyscapes. Furthermore, boundary metrics (BIoU and BF1) show significant improvements with BoundMatch, indicating that our method is more effective in capturing boundary details, while BoundaryMatch has similar boundary metrics to UniMatch.
@@ 4.3.6 Effect of Harmonious BN update strategy
Table 9: Effect of Harmonious Batch-Norm (HBN) update.
Methods VOC (92) Cityscapes ($\nicefrac{{1}}{{16}}$)
IoU $\Delta$ IoU $\Delta$
@@
SAMTH (+ HBN) $70.7$ +7.9 $75.1$ +5.7
Figure 8: Training curves of AugSeg, CutMix-MT, and SAMTH. The bold lines are method trained with HBN update strategy, while the dotted lines are the ones without.
To further examine the effect of our Harmonious Batch Normalization (HBN) update strategy, we compare different BN update approaches in Tab. 9. Methods like CutMix-MT, ReCo, and U2PL also update the teacher’s BN statistics via forward passes. However, unlike HBN (which processes the complete batch through the teacher specifically for BN updates, as described in Sec. 3.3), these methods often introduce discrepancies between the inputs used for the teacher’s BN statistics and the student’s training inputs. Other approaches, such as iMAS and AugSeg, update the teacher’s BN buffers via EMA based on the student’s statistics. As demonstrated in the table, incorporating our HBN strategy consistently yields significant improvements in segmentation performance across various base methods. For example, applying HBN to SAMTH results in a 5.7% absolute improvement in mean IoU on Cityscapes (from 69.4% to 75.1%). The training curves in Fig. 8 provide further evidence, indicating that not only does HBN results in higher final performance, but it also stabilizes the training process.
HBN is not the main focus of this work, but it is a simple yet effective strategy that can be applied to any SS-SS method that uses a teacher-student framework with EMA updates. It led SAMTH, a simple SDA method, to achieve competitive performance on the benchmarks.
@@ 4.3.7 Computational Cost
Table 10: Computation cost comparison. We report the average iteration time in seconds and the GPU memory required by a single GPU during training. We also report FLOPs, number of parameters, and FPS during inference using NVIDIA RTX 3090 GPU.
Method Accuracy (%) Training cost Inference cost
IoU BIoU BF1 Time Mem FLOPs Params FPS
UniMatch $75.3$ $55.1$ $61.0$ $1.21$s $20.6$GB $191$G $40.5$M $23.4$
SAMTH $75.1$ $54.8$ $59.3$ $0.92$s $9.0$GB $191$G $40.5$M $23.4$
+ BoundMatch $76.5$ $56.1$ $62.7$ $1.59$s $16.5$GB $192$G $40.6$M $19.0$
Tab. 10 summarizes the computational costs of our approach during both training and inference. During training, SAMTH+BoundMatch requires more resources than SAMTH: iteration time increases from 0.92s to 1.59s and memory usage rises from 9.0GB to 16.5GB. The increased iteration time primarily stems from the on-the-fly boundary generation pipeline [11], which computes ground-truth boundaries from labeled data at each iteration. The higher memory usage results from processing spatially large feature maps in the boundary detection head, particularly from early backbone stages. Notably, despite these increases, our total memory footprint remains lower than UniMatch (20.6GB) while iteration time gap reduces to only 0.38s.
For inference, the computational overhead is minimal, with only a modest decrease in FPS (from 23.4 to 19.0) while FLOPs and parameters remain nearly unchanged. These costs are justified by the significant performance gains: +1.4% mIoU overall and up to +3.4% improvement in boundary-specific metrics. The training overhead could be reduced through optimizations of the preprocessing pipeline, which we leave for future work. Given that training is a one-time cost while inference benefits are permanent, this trade-off is particularly favorable for applications requiring precise boundary delineation.
@@ 4.4 More Real-World SS-SS Setting
Inspired by the evaluation protocol in [49], we assess SS-SS in a realistic scenario: using all 2,975 Cityscapes training images as labeled data and the additional 19,997 “extra” images as truly unlabeled data. In this setting, the fully supervised baseline is already strong, making further gains challenging yet highly relevant for practical deployment. As shown in Tab. 11, SAMTH+BoundMatch achieves consistent improvements over both SAMTH and UniMatch, particularly in boundary-specific metrics (BIoU and BF1), demonstrating its effectiveness at leveraging large unlabeled pools in real-world applications.
@@ 4.5 Application to lightweight models
Method Accuracy (%) Efficiency
IoU BIoU BF1 FLOPs Params FPS
DLV3+ MV2 Baseline $73.4$ $53.9$ $56.5$ $89.7$G $5.8$M $36.0$
@@
We further evaluate BoundMatch on lightweight segmentation architectures suitable for resource-constrained devices. Specifically, we apply our method to a MobileNet-V2 backbone [64] with DeepLabV3+ and to AFFormer-Tiny, an efficient Transformer model [22]. For the latter, we compare against Mobile-Seed [12], a strong baseline for boundary-aware lightweight segmentation using AFFormer-Tiny.
@@
Notably, when applied to AFFormer-Tiny, BoundMatch not only surpasses the segmentation quality of the specialized Mobile-Seed method but also achieves higher throughput (FPS) with slightly lower FLOPs. This highlights the efficiency benefits of our MTL architecture combined with SS-SS training on additional unlabeled data, even compared to mobile segmentation methods.